40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 12 02:09:39 2014, ericq, Update, General, Reasonable alignment restored 
    Reply  Tue Aug 12 14:29:01 2014, ericq, Update, General, Reasonable alignment restored OLtrend.png
Message ID: 10367     Entry time: Tue Aug 12 02:09:39 2014     Reply to this: 10369
Author: ericq 
Type: Update 
Category: General 
Subject: Reasonable alignment restored 

I took over the IFO, after Jenne's locking efforts, which included manual alignment, since the ASS was doing bad things. 

For whatever reason, the Yarm ASS TT gains needed to be flipped back to go in the right direction. I've restored the old BURT snap file, and the ASS seems to work for now.  

Furthermore, I added some FMs to the Yarm ASS to be able to ramp down gains, to be done as new offsets are ramped in, so that a smooth offset transition is possible. The new version of the script works reasonably, but could be smoother still... Once I iron this out, I'll do the same change to the Xarm, and update the buttons. 

In any case, I was able to run ASS on both arms; single arm lock maxed out at around 0.85, maybe because we're only getting 0.78 from the PMC and 16k from the MC? I then aligned and locked the PRM, then reentered the oplevs on all of the PRMI optics. Oddly, the ETMs were at single uRads on their oplevs.

With this arm alignment, I was able to get the green TRX to ~0.55, and thus the beatnote to around -25dBm, which is still lower than we'd like. I didn't touch the Y green alignment, though it is pretty bad, at transmission of below 0.2 when "locked" on the 00 mode. 

When I try to lock things, the initial ALS CARM and DARM locking seems to go fine, actuating on the ETMs for both DoFs, but ETMX is getting kicked during the resonance search every time. Maybe improving green alignment / increasing beatnote amplitudes will hopefully help some.

I'm leaving the interferometer with the PRM aligned, so that all optics (except SRM) are near the center of their oplev range. I'm curious as to what their variance will be over the next day; this can inform whether we need to improve the ETMY oplev's angular range or not. 

ELOG V3.1.3-